# A decoder-only foundation model for time-series forecasting

### v1

@@ 5.1 Zero-shot Evaluation
$\mathrm{WAPE}\mathopen{}\mathclose{{}\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}\right)$ $=\frac{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}-\hat{\mathbf{y}}_{L+1:L+H}}}\right\|_{1}}{\frac{1}{H}\mathopen{}\mathclose{{}\left\|{\mathbf{y}_{L+1:L+H}}}\right\|_{1}}.$
The metrics in the table are across all time-series in the test period of the target datasets. The metrics are calculated over all rolling window (context, horizon) pairs that can be extracted from the test period. Note that FEDformer training was not able to scale to the Wiki dataset (the largest of our target datasets) and hence its metrics for Wiki in the table are left blank.
For each column in the table (corresponding to a metric computed for a dataset), we color-code the best performance by blue, the second-best performance by green, and the worst performance by red. We observe that PreDcT(ZS) obtained uniformly good results (in the ballpark of the best supervised models) for a majority of the target datasets. In particular, PreDcT(ZS) obtained the best performance for the TourismL dataset, and close to the best performing models for Wiki, ETTh1 and ETTh2. It was among the best or second-best performing model for $8$ out of the $14$ columns in the table, and the worst-performing model for only one of the columns(NRMSE on ILI). This is particularly remarkable since we use a single, pretrained model evaluated in zero-shot manner on the target datasets, and are comparing here against state-of-the-art supervised baselines trained separately on each of the datasets.
@@
! Model ETTh1 ETTh2 ETTm1 ETTm2 Wiki ILI TourismL NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PatchTST(S) 0.656 0.380 0.245 0.161 0.571 0.307 0.180 0.114 0.115 0.081 0.414 0.132 0.590 0.203 TiDE(S) 0.663 0.374 0.245 0.161 0.588 0.320 0.186 0.120 0.103 0.070 0.455 0.131 0.574 0.194 N-BEATS(S) 0.687 0.421 0.235 0.152 0.608 0.320 0.179 0.111 0.103 0.069 0.410 0.137 0.666 0.213 FEDFormer(S) 0.675 0.411 0.294 0.205 0.635 0.381 0.240 0.152 - - 0.441 0.164 1.080 0.266 PreDcT(ZS) 0.672 0.378 0.238 0.151 0.591 0.310 0.199 0.123 0.107 0.070 0.477 0.152 0.514 0.192
Table 2: We present NRMSE and WAPE metrics for (S)upervised models and our (Z)ero-(S)hot model. The supervised models are trained, tuned and evaluated on the specific target datasets. PreDcT(ZS) metrics are reported on zero-shot performance i.e the model has never seen the target dataset prior to inference. The best number in each column is colored blue, and the second-best number is colored green. The worst performing model metric per column is colored red. It can be seen that the PreDcT(ZS) metrics are uniformly good across all datasets and was among the best or second-best performing model for $8$ out of the $14$ columns. We report standard errors for the supervised metrics in Table 7 in the appendix.
5.2 Ablation
Next, we perform several ablation studies that inform the design decisions we made for our foundation model architecture.
Different architectures on same pretraining data. [13] have shown that PatchTST can be used to learn semi-supervised representations of time-series. Similarly  [15] have shown that the N-BEATS architecture can be used for transfer learning in time-series. Therefore, we also consider pre-training a foundation model based on the PatchTST and N-BEATS architecture using the same pretraining dataset, and evaluate them in a zero-shot manner on the target datasets, similar to PreDcT(ZS). These baselines will be denoted by PatchTST(ZS) and N-BEATS(ZS). Note that N-BEATS(ZS) was restricted to training and inference with a fixed context length on account of being a MLP model.
The results are shown in Table 3. It can be seen that PreDcT(ZS) performs better than or similar to PatchTST(ZS) on ETTh1, ETTh2, ETTm2, Wiki and TourismL, but we are dramatically better than PatchTST(ZS) on ETTm2 and ILI. Note that because of encoder-decoder only training PatchTST can only adapt to context lengths that are used for pretraining which are 512, 256 and 64 as mentioned in the Pretraining Data section above. This is evident by its bad performance on the ILI dataset which has a context length of 96. This can be further seen in the study in Table 4, discussed subsequently. N-BEATS(ZS) performs slightly better than us on ETTm2, slightly worse on ETTm1, and similar on ETTh1 and ETTh2. But it cannot adapt to varying context lengths, so it could not generalize to the other datasets for Wiki, ILI and TourismL.
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 Wiki ILI TourismL NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PreDcT(ZS) 0.672 0.378 0.238 0.151 0.591 0.310 0.199 0.123 0.107 0.070 0.477 0.152 0.514 0.192 PatchTST(ZS) 0.670 0.374 0.239 0.151 0.740 0.370 0.205 0.127 0.109 0.073 0.618 0.199 0.505 0.186 N-BEATS(ZS) 0.670 0.381 0.239 0.153 0.617 0.323 0.189 0.119 - - - - - -
Table 3: We present metrics for the three different zero-shot model architectures. It can be see that PreDcT(ZS) does uniformly well across all datasets. PatchTST(ZS) does not do well on ILI on account of not being able to generalize to context 96 because of encoder-decoder mode of training. N-BEATS numbers could not be obtained on non-ETT datasets because it has a fixed context length due to its MLP architecture. The best number in each column is made bold.
Adapting to different context lengths.  A good foundation model should be able to adapt to a variety of different context lengths. This is possible in our model because of decoder-only training - the output token of every patch extracts features from all the patches that come before it, and is trained to predict the next output patch. In Table 4 we show the performance (in terms of NRMSE) of PreDcT(ZS) with different context lengths on the same task as before of predicting 96 time-points. We also juxtapose our performance with PatchTST(ZS) that is trained in encoder-decoder fashion. It can be seen that our model has good performance throughout which becomes progressively better with more context. On the other hand the performance of PatchTST is only good for context length 512 because it has not been optimized for other context lengths because of encoder-decoder model training. Note that because of overlapping stride the original PatchTST model does not lend itself easily to decoder-only training.
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 96 192 384 512 96 192 384 512 96 192 384 512 96 192 384 512 PreDcT(ZS) 0.779 0.715 0.692 0.672 0.253 0.250 0.239 0.238 0.663 0.631 0.607 0.591 0.204 0.202 0.201 0.199 PatchTST(ZS) 1.429 1.125 0.995 0.670 0.333 0.279 0.245 0.239 1.168 1.327 1.106 0.740 0.320 0.261 0.220 0.205
Table 4: NRMSE numbers are presented for the pretrained models when the context length is varied at inference time. The prediction horizon is held fixed at 96. It can be seen that PreDcT(ZS) can adapt to different context lengths at inference time.
Input patch length.  The size of input_patch_len represents an important trade-off. We have typically seen that increasing its value from 8 to 64 increases performance but having too high a input_patch_len is impractical because the model cannot be easily applied to context lengths that are less than input_patch_len, at inference time. In many monthly and higher granularity tasks, it is common to have small context lengths. In Table 5 we show the NRMSE of another PreDcT(ZS) model with input_patch_len=8 on ETT datasets, which is clearly worse than our original model that uses input_patch_len=32.
input_patch_len ETTh1 ETTh2 ETTm1 ETTm2
32 0.672 0.238 0.591 0.199
8 0.680 0.263 0.706 0.209
Table 5: Ablation with respect to input patch length. NRMSE numbers are reported.
Autoregressive decoding. In recent long-term forecasting works [24, 13, 5] it has been observed that directly predicting the entire forecasting horizon in one shot from a decoder can yield better results than auto-regressive decoding on long horizon benchmarks. For a foundation model the horizon length of the task is not known before inference time, therefore one-shot decoding might not be possible for very long horizons. However, by keeping the output_patch_len longer than input_patch_len one can ensure fewer autoregressive steps. This was one of the key decisions in the design of PreDcT, that is quite different from LLMs. In order to showcase this we choose the task of predicting 512 time-steps into the future for the ETT datasets. In Table 6, we present results from a model with output_patch_len=32 vs our original model that uses output_patch_len=128. The former has to perform 16 autoregressive steps while the latter has to do only 4. It can be clearly seen that having a larger output_patch_len helps in this case.
output_patch_len ETTh1 ETTh2 ETTm1 ETTm2
128 0.724 0.286 0.711 0.256
32 0.746 0.292 0.737 0.259
Table 6: Ablation with respect to output patch length for the task of predicting 512 steps into the future. NRMSE numbers are reported.
6 Discussion and Future Work
We train a decoder-only foundation model for time-series forecasting using a large pretraining corpus of about 1B time-points, the majority of it being search interest time-series derived from Google trends. We show that even a relatively small 225M parameter pretrained model that uses our PreDcT architecture displays impressive zero-shot performance on a variety of public benchmarks from different domains and granularities. The PreDcT(ZS) model can rival the performance of recent state-of-the-art supervised baselines that have been specifically trained on the target datasets. This is remarkable since the PreDcT(ZS) model has not seen the target datasets before inference.
In future work, it would be interesting to push the boundaries of scale in both the pretraining data as well as the number of parameters in the model. It would be a insightful to perform a scaling study in the lines of [7] for time-series foundation model. An empirical analysis of probabilistic zero-shot forecasting from an appropriately trained foundation model is also an interesting direction for future work.
Appendix A Appendix
A.1 Date Derived Features
In our study we focus on 5 date derived features: (1) month of the year, (2) day of the week, (3) hour of the day, (4) minute of the hour and (5) second of the minute. For any time point $t$ in the context and the horizon we have $\mathbf{x}_{t}\in\mathbb{R}^{5}$. In addition for each date derived feature:
  1. 1.
    We mask it by -1 if its granularity is irrelevant to the time series $\mathbf{y}$. For example, the daily Wikipedia makes no use of hour of the day, minute of the hour, and second of the minute.
  2. 2.
    We normalize its values to $[-0.5,0.5]$. For example, the raw minute of the hour value $v$ is transformed to $v/60-0.5$.
A.2 More Details on Models
N-BEATS. All N-BEATS(S) and N-BEATS(ZS) models are trained following the N-BEATS-G hyper-parameters used in Table-18 from the original paper [14], except that we use width = 1024 and stacks = 24 to match the model sizes to about 200M. All supervised N-BEATS(S) use the dataset specific input output lengths, while N-BEATS(ZS) is trained with input length = 512 and output length = 96 to predict on all ETT datasets.
PatchTST. For PatchTST(S) models in all ETT datasets and ILI we use the hyper-parameters from the original paper [13]. For Wiki and TourismL, we use 16 attention heads, 10 layers, a patch-size of 64 and a stride of 32. The model dimension used is 1024. We tune the learning rate per dataset.
For PatchTST(ZS) we use we use 16 attention heads, 20 layers, a patch-size of 32 and a stride of 16. We use the same model dimension as PreDcT(ZS).
TiDE. We used the hyper-parameters from the original paper [5], but tuned the model dimensions (from 256 to 1024), number of layers (from 1 to 4) and learning date per-dataset.
FEDFormer. For all datasets we use the same hyper-parameters of the original paper [25], except model dimensions. We tuned the model dimension per dataset from 64 to 1024.
PreDcT. We use use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128. The model dimension is set to 1280.
A.3 Additional Empirical Results
[…]
! Model ETTh1 ETTh2 ETTm1 ETTm2 NRMSE WAPE NRMSE WAPE NRMSE WAPE NRMSE WAPE PatchTST(S) 0.656 $\pm$ 0.0004 0.380 $\pm$ 0.0002 0.245 $\pm$ 0.0004 0.161 $\pm$ 0.0004 0.571 $\pm$ 0.003 0.307 $\pm$ 0.001 0.180 $\pm$ 0.0008 0.114 $\pm$ 0.0007 TiDE(S) 0.663 $\pm$ 0.0001 0.374 $\pm$ 0.0001 0.245 $\pm$ 0.002 0.161 $\pm$ 0.001 0.588 $\pm$ 0.0005 0.320 $\pm$ 0.0005 0.186 $\pm$ 0.0005 0.120 $\pm$ 0.0001 N-BEATS(S) 0.687 $\pm$ 0.002 0.421 $\pm$ 0.002 0.235 $\pm$ 0.0002 0.152 $\pm$ 0.0001 0.608 $\pm$ 0.007 0.320 $\pm$ 0.004 0.179 $\pm$ 0.001 0.111 $\pm$ 0.0003 FEDFormer(S) 0.675 $\pm$ 0.003 0.411 $\pm$ 0.004 0.294 $\pm$ 0.010 0.205 $\pm$ 0.01 0.635 $\pm$ 0.007 0.381 $\pm$ 0.005 0.240 $\pm$ 0.006 0.152 $\pm$ 0.004 Model Wiki ILI TourismL - NRMSE WAPE NRMSE WAPE NRMSE WAPE - - PatchTST(S) 0.115 $\pm$ 0.0002 0.081 $\pm$ 0.0002 0.414 $\pm$ 0.005 0.132 $\pm$ 0.003 0.595 $\pm$ 0.012 0.204 $\pm$ 0.003 - - TiDE(S) 0.103 $\pm$ 0.0004 0.070 $\pm$ 0.0006 0.455 $\pm$ 0.004 0.131 $\pm$ 0.005 0.574 $\pm$ 0.001 0.194 $\pm$ 0.002 - - N-BEATS(S) 0.103 $\pm$ 0.0002 0.069 $\pm$ 0.0001 0.410 $\pm$ 0.003 0.137 $\pm$ 0.002 0.666 $\pm$ 0.002 0.213 $\pm$ 0.0003 - - FEDFormer(S) - - 0.441 $\pm$ 0.0007 0.164 $\pm$ 0.0003 1.080 $\pm$ 2 0.266 $\pm$ 0.0007 - -
Table 7: NRMSE and WAPE confidence intervals (mean $\pm$ 1 standard error) of the supervised baselines.

### v2

@@ A.1 Metrics
    $\mathrm{msMAPE}\mathopen{}\mathclose{{\left(\mathbf{y}_{L+1:L+H},\hat{\mathbf{y}}_{L+1:L+H}}}\right)\\ =\frac{1}{H}\sum_{i=1}^{H}\frac{2\lvert y_{L+i}-\hat{y}_{L+i}\rvert}{\max\mathopen{}\mathclose{{\left\{\lvert y_{L+i}\rvert+\lvert\hat{y}_{L+i}\rvert+\epsilon,0.5+\epsilon}}\right\}}.$ (8)
    In Monash benchmarks [8] $\epsilon=0.1$ was used. This metric is used in order to avoid undefined values in other normalized metrics like MAPE. In multivariate datasets the metrics are calculated for each time-series and then we take the mean or the median. In this paper we only use the mean versions.
Aggregating across datasets. Since the datasets have wildly different scales averaging unnormalized metrics like MAE is not kosher. Therefore following [9] we scale the metric of each baseline for a dataset by the same metric achieved by a naive baseline on that dataset. The naive baseline just makes the constant prediction $y_{L}$ repeated across the prediction length. We did not need to do that for the Informer datasets since on these datasets metrics are usually reported on standard normalized data [19].
A.2 Additional Empirical Results
In this section, we provide more detailed tables for our zero-shot datasets and experiments described in Section 6.1
A.2.1 Darts
We present the MAE results individually from all 8 datasets in Table 1. It can be seen that TimesFM performs well for all datasets with clear seasonal patterns. We do not perform the among the top-3 only in Sunspot and Wooly datasets. On an average we are within significant level of the best model and a close second on the mean scaled MAE metric.
Table 1: MAE for Darts datasets. We also include the naive baseline that predicts the last values in the context repeatedly.
gp arima TCN N-BEATS N-HiTS llmtime(ZS) TimesFM(ZS) NAIVE
AirPassengersDataset 34.67 24.03 54.96 97.89 59.16 34.37 14.75 81.45
AusBeerDataset 102.05 17.13 30.90 10.39 34.23 16.13 10.25 96.35
GasRateCO2Dataset 2.27 2.37 2.64 2.63 3.85 3.50 2.69 2.29
MonthlyMilkDataset 30.33 37.19 70.86 33.64 32.73 9.68 22.46 85.71
SunspotsDataset 53.74 43.56 51.82 73.15 49.93 47.34 50.88 48.24
WineDataset 4552.06 2306.70 3287.14 4562.02 3909.51 1569.32 2462.11 4075.28
WoolyDataset 649.98 588.78 1158.79 903.01 382.09 808.73 917.10 1210.33
HeartRateDataset 5.65 5.56 5.49 6.57 6.10 6.21 5.44 5.92
A.2.2 Monash
Table 2: We present the mean MAE results for our methods along size monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
Dataset
australian electricity demand 459.96 659.60 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.50 231.45 523.12 659.60
bitcoin 1.75e18 5.33e18 5.33e18 9.9e18 1.10e18 3.62e18 6.66e18 1.93e18 1.45e18 1.95e18 1.06e18 2.46e18 2.61e18 1.97e18 5.32e18
pedestrian counts 70.20 170.87 170.94 222.38 216.50 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29 58.35 170.88
weather 2.32 2.24 2.51 2.30 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03 2.17 2.36
nn5 daily 9.39 6.63 3.80 3.70 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16 4.09 8.26
nn5 weekly 15.91 15.66 15.30 14.98 15.70 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34 14.33 16.71
tourism yearly 140081.78 95579.23 90653.60 94121.08 94818.89 95033.24 82682.97 79567.22 79593.22 71471.29 70951.80 69905.47 74316.52 93572.42 99456.05
tourism quarterly 14121.09 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67 10037.84 15845.10
tourism monthly 4724.94 5302.10 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98 2916.02 5636.83
cif 2016 715086.33 581875.97 714818.58 855578.40 642421.42 469059.49 563205.57 603551.30 1495923.44 3200418.00 679034.80 5998224.62 4057973.04 896414.41 386526.37
covid deaths 304.68 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66 555.24 2653.98
fred md 2013.49 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.80 2508.40 4666.04 1823.16 2825.67
traffic hourly 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.03
traffic weekly 1.17 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.20 1.42 1.13 1.19
saugeenday 28.63 21.50 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06 24.53 21.50
us births 459.43 1192.20 586.93 399.00 419.73 526.33 574.93 441.70 557.87 424.93 422.00 504.40 452.87 408.49 1152.67
hospital 24.62 21.76 18.54 17.43 17.97 19.60 19.24 19.17 22.86 18.25 20.18 19.35 36.19 19.34 24.07
solar weekly 2049.09 1202.39 1210.83 908.65 1131.01 839.88 1044.98 1513.49 1050.84 721.59 1172.64 1996.89 576.35 1184.01 1729.41
Table 3: We present the mean msMAPE results for our methods along size monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
Dataset
australian electricity demand 15.30 22.07 22.18 13.70 44.22 28.75 8.76 8.35 9.22 12.77 7.74 8.07 8.68 17.27 22.07
bitcoin 22.70 30.31 40.36 20.12 20.65 31.11 21.48 29.78 21.00 21.16 31.95 22.16 23.32 42.00 192.46
pedestrian counts 53.02 121.39 122.08 119.76 148.48 138.58 40.29 45.54 39.30 36.10 54.15 34.22 36.16 50.73 121.35
weather 34.05 50.85 56.19 58.06 51.47 57.98 106.01 59.12 38.17 36.46 50.87 40.13 35.35 50.85 36.81
nn5 daily 47.43 35.38 21.93 21.11 21.49 25.91 30.20 24.04 23.30 23.75 28.47 22.65 23.18 23.46 48.09
nn5 weekly 12.28 12.24 11.96 11.62 12.29 11.83 11.45 11.67 11.49 11.52 10.93 14.95 14.83 11.01 13.26
tourism yearly 48.28 34.10 31.93 33.94 36.52 33.39 46.92 31.54 33.73 34.06 30.24 28.80 34.66 30.14 42.14
tourism quarterly 25.25 27.41 15.37 17.16 15.07 16.58 15.86 16.53 16.20 15.29 14.45 15.56 16.97 19.92 31.68
tourism monthly 33.53 36.39 19.89 21.20 19.02 19.73 21.11 21.10 20.11 18.35 20.42 18.92 19.74 22.77 40.40
cif 2016 18.63 14.94 13.04 12.19 12.18 11.69 12.32 14.86 12.31 13.54 11.71 18.82 12.55 20.00 15.41
covid deaths 20.65 15.35 15.57 8.71 8.64 9.26 18.34 15.40 18.58 34.18 32.36 14.50 40.71 15.89 134.79
fred md 9.75 8.72 9.72 7.97 8.40 7.98 30.77 9.16 8.99 8.33 8.32 9.08 11.26 8.47 8.76
traffic hourly 11.70 8.73 8.73 12.58 9.84 11.72 5.97 7.94 4.30 4.16 5.16 5.22 4.10 4.76 8.73
traffic weekly 12.85 12.40 12.48 12.80 12.63 13.45 12.46 12.90 12.64 13.13 12.31 13.21 15.18 12.48 12.98
saugeenday 58.53 35.99 35.97 37.34 67.50 37.55 45.32 35.55 39.32 40.22 56.03 37.02 56.62 43.31 35.99
us births 4.48 11.77 5.82 3.81 4.05 5.17 5.75 4.23 5.55 4.13 4.17 4.88 4.36 3.94 11.35
hospital 21.68 17.94 17.27 17.55 17.46 17.79 17.56 18.04 18.29 17.41 17.72 17.51 20.04 17.86 21.55
solar weekly 37.33 24.59 24.76 19.05 22.93 17.87 21.65 29.35 21.52 15.00 24.05 32.50 12.26 24.20 32.80
In Table 2 we present the actual MAE numbers that are behind the main Figure 2(a).
It should be noted that though the scaled MAE metric accounts for scale variation across datasets, it does not account for within-dataset scale variations. Indeed, in datasets like cif 2016 and covid deaths, there are time-series with different order of magnitudes. The mean MAE for these datasets does not reflect average performance but only the performance of the higher scale time-series. Therefore in Table 3 we also present the normalized msMAPE metrics for TimesFM and all baselines.
A.2.3 Informer
We present the MAE on the last split of the test set for all dataset, horizon pairs considered in Table 4. Owing to expensive evaluations for llmtime, the results are reported on the last test window of the original test split, as done in [9].
Table 4: MAE for ETT datasets for prediction horizons 96 and 192. Owing to expensive evaluations for llmtime, the results are reported on the last test window of the original test split.
llmtime(ZS) PatchTST TimesFM(ZS) FEDFormer AutoFormer Informer
Dataset
ETTh1 (Horizon=96) 0.42 0.41 0.37 0.58 0.55 0.76
ETTh1 (Horizon=192) 0.50 0.49 0.49 0.64 0.64 0.78
ETTh2 (Horizon=96) 0.33 0.28 0.28 0.67 0.65 1.94
ETTh2 (Horizon=192) 0.70 0.68 0.58 0.82 0.82 2.02
ETTm1 (Horizon=96) 0.37 0.33 0.25 0.41 0.54 0.71
ETTm1 (Horizon=192) 0.71 0.31 0.24 0.49 0.46 0.68
ETTm2 (Horizon=96) 0.29 0.23 0.28 0.36 0.29 0.48
ETTm2 (Horizon=192) 0.31 0.25 0.24 0.25 0.30 0.51
A.3 More Details on Models
We now present implementation details about TimesFM and other baselines.
Monash Baselines. The raw metrics for the Monash baselines are directly taken from Tables 9 and 11 of the supplementary material of the original paper [8]. For llmtime, we use the precomputed outputs provided by the authors of  [9].
Darts Baselines. For all the Darts baselines we use the precomputed outputs provided by the authors of  [9]. For more details please see Section C.1 in that paper.
TimesFM. For our main 200M model we use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128. The model dimension is set to 1280. We train with layer norm and a cosine decay learning rate schedule with peak learning rate of $5e-4$.
Informer Baselines. For FEDFormer [35], Autoformer [32], Informer [34] and PatchTST [19] we use the original hyperparameters and implementation. The results presented in the main paper are obtained on the last test window of length horizon length as stated in the llmtime [9] paper.
We generate the llmtime predictions using the code provided by the authors  [^6] https://github.com/ngruver/llmtime/blob/main/experiments/run_monash.py but adapted to the ETT datasets. Note that as of January 2024, OpenAI has discontinued access to GPT-3, therefore we had to use the GPT-3.5-Turbo model.
A.4 Synthetic Data
We create the synthetic data to reflect common time series patterns using traditional statistical models. We start with four simple times series patterns:
  • Piece-wise linear trends (I), where the number of the piece-wise linear components is randomly chosen between 2 and 8.
@@ A.4 Synthetic Data
  • Seasonal patterns. In particular we create the sine (III) and the cosine (IV) waves of different random periods between 4 and 96 time-points and time delays.
We then randomly enable / disable these four components (I) - (IV), generate their time series of length 1024 respectively, and sum them up using uniformly sampled random weights to create each times series in the synthetic datasets. We also choose to apply the trend multiplicatively 50% of the times the trend component is chosen.